OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20 Please help!

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \sf |2x + 6| − 4 = 20 }\) add 4 to both sides \(\Large\color{blue}{ \sf |2x + 6| = 16 }\) now you have 2 equations, \(\Large\color{red}{ \sf 2x + 6 =16 }\) and \(\Large\color{red}{ \sf 2x + 6 =-16 }\) can yu solve each of the red equations for x, show your steps and the answer please.

OpenStudy (anonymous):

2x+6=16 -6 -6 2x=10 divide both sides by two i got X=5 2x+6=-16 -6 -6 2x= -22 divide by two x= -11

OpenStudy (anonymous):

The answer options are x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No solution I don't know how they got it

OpenStudy (solomonzelman):

OHH, SORRY !

OpenStudy (solomonzelman):

I didn't add 4 to each side correctly. @NatPanda see what Luigi posted? try to apply this formula to your example. I'll start you of. \(\Huge\color{blue}{ \sf |2x + 6| − 4 = 20 }\) \(\Huge\color{blue}{ \sf |2x + 6| = 24 }\) (24 !! not 16) \(\Large\color{green}{ \sf 2x + 6 = 24 }\) and \(\Large\color{green}{ \sf 2x + 6 =24 }\)

OpenStudy (solomonzelman):

The second green equation is -24

OpenStudy (anonymous):

2x+6=24 -6 -6 2x=8 x=9 2x+6=-24 -6 -6 2x= -30 x= - 15

OpenStudy (anonymous):

Thank you!

OpenStudy (solomonzelman):

Hold on

OpenStudy (solomonzelman):

you had 1 mistake here \(\Large\color{green}{ \sf 2x + 6 = 24 }\) \(\Large\color{green}{ \sf 2x = }\)\(\Large\color{red}{ \sf 18}\) second one, x=-15 is correct.

OpenStudy (anonymous):

i got 18 i accidentally forgot to type the one

OpenStudy (solomonzelman):

Oh, I see, but the answers are all correct. Good work !

OpenStudy (anonymous):

Thank you!

OpenStudy (solomonzelman):

My pleasure !